Jose left the airport and traveled toward the mountains. Kayla left 2.1 hours later traveling 35 mph faster in an effort to catch up to him. After 1.2 hours Kayla finally caught up. Find Jose's average speed.
Answer by stanbon(75887)

Jose left the airport and traveled toward the mountains. Kayla left 2.1 hours later traveling 35 mph faster in an effort to catch up to him. After 1.2 hours Kayla finally caught up. Find Jose's average speed
------
Jose DATA:
time = 2.1+1.2 = 3.3 hrs ; rate = x mph ; distance = 3.3x miles
---------------
Kayla DATA:
time = 1.2 hrs ; rate = x+35 mph ; distance = 1.2(x+35) = 1.2x + 42 miles
---------------------
Equation:
distance = distance
3.3x = 1.2x + 42
2.1x = 42
x = 20 mph (Jose's average speed)
=======================================
Cheers,
Stan H.